Assuming a pre-thermonuclear test equilibrium value of 1.3 TU (tritium units) for tritium in a specific grape-growing region, calculate the age of a wine produced there that has a 3H/1H of 8.35 x10-20. 1 TU = 1 atom 3H/1 x 1018 atoms of 1H. The half-life of tritium is 12.33 years.
Answer #1

equilibrium value of 3H/1H is X_0=1.3\,TU=\frac{1.3}{1\times10^{18}}

value of 3H/1H in wine is X = 8.35\times10^{-20}

Half-life of Tritium T_{1/2}=12.33\,years

since X=X_0e^{-\lambda{t}}=X_02^{-\frac{t}{T_{1/2}}}

8.35\times10^{-20}=\frac{1.3}{1\times10^{18}}2^{-\frac{t}{T_{1/2}}}

2^{-\frac{t}{T_{1/2}}}=\frac{8.35\times10^{-2}}{1.3}

2^{\frac{t}{T_{1/2}}}=\frac{1.3\times10^{2}}{8.35}=\frac{130}{8.5}

t=48.52\,years